A light rail commuter train decelerates at a rate of 1.65m/s2. If it’s traveling at a max speed of 80km/h, how long does it take to come to a stop? Round two decimal places.


1
Expert's answer
2021-02-15T17:38:19-0500
"t=\frac{v}{a}\\t=\frac{80}{1.65\cdot 3.6}=13.47\ s"